OpenStudy (anonymous):

How long will it take a car to go 200 km traveling at v- 5 km/h

OpenStudy (anonymous):

V = D/T

OpenStudy (anonymous):

wait but how would you solve? can you plug in the numbers and show me how to get to the answer?

OpenStudy (anonymous):

t = d/v t = 200/5 t = 40 h